1. A bullet train accelerates from rest at Tokyo Station to 89 m/s over a distance of 120 m. What acceleration did it experience?
1
Expert's answer
2022-01-04T16:17:35-0500

The acceleration is


"a=v^2/(2d)=89^2/(2·240)=16.5\text{ m/s}^2."